# Video Is Worth a Thousand Images: Exploring the Latest Trends in Long Video Generation

### v1

@@ 3.2.1. Large Language Model As Director
The Large Language Model as Director  [55] [125]  [86],[86] [38] paradigm starts by identifying key frames that define the core narrative and then creates the intermediate frames to smoothly connect these critical frames into a cohesive, extended video. This approach differentiates between essential storyline keyframes and the additional frames required for continuity. It consists of two main components: the LLM Planner, which produces the storyline blueprint, and the Video Generator Backbone, which follows the blueprint provided by the LLM to generate the video as illustrated in Fig. 9. Initially, the model develops a script detailing each frame and optional metadata, such as entities, layouts, and actions. The model can then create images and interpolate between them using the semantic descriptions and entities from earlier frames to guide the denoising process. LLMs are crucial in this approach, as they generate detailed storylines from text prompts, specifying scenes, characters, and actions  [55]. This paradigm works with both training-based and training-free (zero-shot) approaches. Free-Bloom  [37] took a zero-shot approach. It has a three-stage architecture starting with LLM to generate a sequence of text-based prompts that decompose the original prompt into a semantically coherent sequence of prompts using world knowledge of LLM. It proposes novel techniques like joint noise sampling, step-aware attention, and dual-path interpolation to ensure the temporal coherence of the generated frames. The model implements joint noise sampling by combining unified noise applied at the video level with individual noise applied at the frame level. By making the attention mechanism aware of each specific frame in the generation process, the model dynamically adjusts its focus on different aspects of the input data, including the text prompt and previously generated frames. In dual interpolation, the contextual path interpolates the latent variables between the contextual frames to ensure temporal coherence. The denoising path interpolates latent variables by DDIM  [93] denoising process conditioned on interpolated text embedding to improve semantic coherence. One of the limitations of Free Bloom is its limited use of LLM script generation capacity, as LLM can create a rich script with storylines, scene descriptions, and entities. Video Director GPT  [55] is built in a zero-shot setting, similar to Free Bloom, but it provides much more detailed guidance from GPT-4 as a video planner. It develops a comprehensive story blueprint with multiple components, including layouts, bounding boxes, and entities like actors and objects. This plan is executed using Layout2Vid, based on ModelScopeT2V [125], a grounded video generation module that ensures layout and consistency control across multi-scene videos. The LLM-grounded video diffusion model  [53], similar to VideoDirectorGPT, uses detailed guidance from LLM, such as frame layout and trajectory information. However, it is a training-based approach instead of a zero-shot approach. It can understand complex spatiotemporal dynamics from text and generate layouts that align closely with the prompts and the object motion patterns typically observed in the real world.
Figure 8. LLM as director approach uses LLM as the spatiotemporal director of the script and a separate video generation module that can understand the DSL(metadata) generated by LLM.  [52]
@@
Like VideoDirectorGPT, FlowZero  [62] took a zero-shot (training free) approach, but LLM plays a more detailed role than Video Director GPT. It generates a detailed dynamic scene syntax (DSS), including scene descriptions, object arrangements, and background motion patterns. The DSS components direct an image diffusion model to generate videos with smooth object movements and consistent frame transitions. The theme of using dynamic scene layout is illustrated in Fig. 8. The LLM as director approach has limitations as it is a two-stage architecture, and adding speech or integrating short clips will need modification in the pipeline. We can extend the LLM-based divide-and-conquer approach by incorporating more specialized components, such as a model for generating reference images, a module for video creation, and a plug-and-play module for adding transitional clips and speech. That is achieved using the multi-agent framework, as described in the section 3.2.2.
@@ 3.2.2. Multi stage/Agent Based Divide And Conquer Approach
Figure 10. Mora utilizes a multi-agent framework. The prompt selection agent enhances prompts with detailed instructions, the text-to-image image agent generates images from the input prompts, the image-to-image agent enhances the quality of the photos, the text-to-video agent generates video segments, and the video transition agent integrates these videos into a longer video  [145].
An Agent-Based LLM Framework  [146] is a system where multiple models function as specialized agents, each designed to perform a distinct task or interact within a defined environment. In this system, the LLM serves as the ”brain,” responsible for overseeing complex operations, while simpler models function as tools, executing more specific, supportive tasks. A multi-agent framework for video generation represents a multi-layered approach to text-to-video generation. It produces high-quality long videos like those generated by Sora  [58] by dividing the video creation challenge into multiple systems, each specializing in some aspect of the video generation pipeline as illustrated in Fig. 10. The framework facilitates efficient and dynamic video production by coordinating these agents flexibly and iteratively. Systems like Mora [145] and Vlogger [158] illustrate how integrating and coordinating multiple specialized agents can enhance and streamline the video creation. In Mora [145] Multiple models collaborate through an agent framework for script creation, image generation, image enhancement, video production, and video patching. VLogger  [158] is a comprehensive system designed to produce cohesive vlogs by integrating various modules. It begins with the LLM Director translating a user’s story into a detailed script outlining scenes and duration. The process continues with the Actor module, where the LLM Director assesses the script to identify characters and collaborates with a character designer like SD-XL [74] to generate appropriate reference images for the actors. It has a show maker module to act as a videographer and a voicer module that synchronizes the script with voice-over, utilizing a model like Bark  [111] to align subtitles with the video. This setup exemplifies a modular agent-based system for high-quality video production. DreamFactory  [135] simulates an AI-driven film production team, with LLM-based agents assuming roles such as directors, art directors, screenwriters, and artists. These agents collaborate on script writing, storyboarding, character design, keyframe creation, and video synthesis. DreamFactory takes an approach similar to Mora. Still, it has enhanced the script creation part by using multiple LLM models as different movie actors, like a screenwriter and director, and these agents engage in iterative conversions, finalizing scripts and action plans through the Chain of Thought (COT) pattern. Kubrick  [26] employs a similar collaborative approach for script generation involving LLM agents in dialogue to finalize scripts based on user prompts. It divides the video generation into phases managed by the Director, Programmer, and Reviewer agents. Utilizing iterative reasoning, Kubrick leverages the Blender engine [100] to produce realistic 3D animations. Another way the divide-and-conquer approach can be used is to view a long video as a composition of short videos, first generating these short videos and then integrating them into a long video. That will be discussed in the next section.
@@ 3.2.3. Divide and Conquer Compositional/Transition Approach
This method uses or generates short video segments using any standard video generation technique, such as divide-and-conquer strategies or autoregressive models. It then integrates these segments by employing a transition model to create smooth transitions between them. The challenge with transition videos is to ensure coherence and visual quality and smoothly transition from one set of entities to another. SEINE  [11] adopts a predictive approach for creating transitions between short videos produced by any public text-to-video model. It uses the last frame of one segment and the first frame of the next as input, applying a random masked diffusion model to handle the transitions. This approach involves jointly denoising overlapping short videos in the temporal domain. It starts by generating these short videos with a standard diffusion model, incorporating temporal overlaps based on given prompts. The model then stitches these segments together into a longer video. The process includes transforming the video into a reduced-noise domain through noise inversion and additional denoising with the GenL model. MEVG  [68] ensures temporal coherence between independently generated video clips based on multiple event descriptions. It first creates a video for each event description using a public single-prompt video generation model. Subsequent clips are then generated by conditioning on the last frame of the previous clip and the provided text, maintaining continuity. MAVIN  [147](Multi-Action Video INfilling model) is designed to generate transition videos that seamlessly connect two given videos, forming a cohesive integrated sequence. It learns video infilling by dividing a training video into three segments, corrupting the intermediate segment, and then learning to predict the noise of the missing segment. ENCODER-EMPOWERED GAN [140] generates long videos using an encoder-based GAN to connect short generated video clips into longer sequences of hundreds of frames. It models temporal relationships between consecutive video clips. The process involves generating short video clips with EncGAN3 and then training to enforce temporal connectivity between these clips through a recall mechanism.
The autoregressive approach ensures smooth transitions between frames by generating each frame conditioned on the previous ones, but its sequential nature makes it inherently slow for long video generation. The divide-and-conquer 3.2.1 approach can generate keyframes in parallel but suffers from challenges involving interpolation between frames with smooth transitions and higher video quality while maintaining semantic consistency. Implicit generation 3.3 approaches combine the best of both worlds by generating complete videos directly from a model conditioned on user input without the need for interpolation (divide and conquer) or extrapolation (autoregressive) between frames. A brief summary of some key papers that explore themes related to the Divide and Conquer approach is provided in Table  LABEL:table:divideandconquercatalog.
@@
SEINE [11] Long video with smooth transitions from short vidoes. Oct 2023 integration
@@
3.3. Implicit Video Generation
Implicit video generation covers a family of models that generate complete video using compact latent space representation, enhanced attention mechanisms to incorporate long-term dependencies, or denoising strategies to include short-term and global noise in the generated video. What differentiates implicit video generation from divide-and-conquer and the autoregressive approach is that it generates all frames simultaneously instead of sequentially or parallelly. For example, Sora  [57] compresses raw video into a latent spacetime representation, extracting patches to capture visual appearance and motion dynamics over short intervals. As a diffusion transformer, Sora consists of three components: (1) a time-space compressor that maps the original video into latent space, (2) a Vision Transformer (ViT) that processes this latent representation to produce a denoised version, and (3) a CLIP-like conditioning mechanism that uses LLM-augmented instructions and visual prompts to guide video generation. This is illustrated in 11. FreeNoise  [75] presents a tuning-free and time-efficient paradigm that enhances pre-trained video diffusion models’ generative capabilities while maintaining semantic consistency. GLOBER  [97] is a video auto-encoder that includes a video encoder to compress videos into global features and a video decoder based on a diffusion model to reconstruct video frames from these features in a non-autoregressive manner.
Despite progress in these areas, motion consistency, semantic alignment, and parallel processing remain key obstacles to achieving scalable, high-quality long video generation. An example of this limitation can be seen in state-of-the-art video generation models like SORA [110]. For instance, SORA generates videos that are only up to one minute long, far shorter than the typical duration of entertainment or educational videos. Issues such as unrealistic and incoherent motion, intermittent object appearances, and unrealistic phenomena continue to plague its output, as documented by  [98]. Some key papers that explore themes related to the implicit approaches are cataloged in Table  3.
@@ 3.3. Implicit Video Generation
In addition to generation strategy, another theme in the long videos is input control mechanisms like text, bounding boxes, and images for video guidance. We will discuss that in the next section.
@@ 4. Long Video: Input Control
Poplar mechanisms for input conditioning of long videos are User Textual Prompt, User Textual Prompt With Scene Layout, and Image Input With Textual Prompt And Scene Layout, which will be discussed next.
@@ 4.1. User Textual Prompt
User textual prompts are the most common way of conditioning the video generation models. Video generation can utilize single or multiple text prompts. Dall-E  [80] is one of the first transformer-based text-to-image generation models. Dall-E concatenated BPE-encoded text tokens with the image tokens and trained an autoregressive transformer to model the joint distribution over the text and image tokens. CogVideo[34] takes a similar approach and trains autoregressive transformers with text and image tokens together on joint distribution. CogVideo has conditioning limitations to a single prompt. Phenaki  [121] improved this paradigm by incorporating t5x  [82] pre-trained embedding and conditioned video generation on the sequence of prompts in story fashion. During inference in the autoregressive phase, the model conditions the next frame on the previous frame and text embeddings, which can either remain the same or be based on a new prompt. Phenaki has incoherent motion and semantic transition between video segments from one prompt to another, as these sequential prompts need more semantic coherence. Different approaches attempt this challenge by generating long videos and managing multiple prompts while ensuring temporal consistency and alignment with detailed frame-level descriptions, often with the help of LLM. Free-Bloom  [37] addresses this issue by using LLM to generate semantic coherent prompts combining global and local noise distributions for each frame and employing spatial-temporal self-attention to integrate current and previous features. LLM Grounded Video Diffusion  [52] utilize a multi-modal approach that alternates between steps guided by domain-specific language detailing entities, backgrounds, and scene layouts and denoising steps within the video diffusion model. VideoDrafter [60] introduces modifications to cross-attention maps in image and video generation diffusion models, enabling the production of images and videos that align with both input prompts and reference images, highlighting changes made to the cross-attention mechanism. DirecT2V  [33] utilized LLM like GPT4 to generate step-by-step textual prompt instructions.
@@
Fig. 12 illustrates the DirecT2V  [33] architecture, which modifies the attention block of U-Net and incorporates modulated self-attention.Text-only prompts can guide long video generation but lack semantics for fine-grain control over long video generation. In addition to frame description, adding metadata like bounding boxes of entities like persons and cars and background can help generate a more accurate depiction of videos and help with fine alignment between text and video.
@@ 4.2. User Textual Prompt with Scenes layout
Multi-modal input control mechanisms can work with multi-modalities like text, images, edge maps, bounding boxes, music, etc. We can enhance text-only prompts by adding metadata about bounding boxes, entity descriptions, trajectories, and background context, as illustrated in Fig. 8. Stable Diffusion did some groundwork for multi-modal input control  [84]. It conditions diffusion-based general conditioning mechanisms based on cross-attention by augmenting their underlying UNet backbone as illustarted in Fig. 13. ControlNet  [150] lays out a groundbreaking approach to managing multi-modal control diffusion. It fine-tunes an existing diffusion-based network by leveraging a large pre-trained model as a robust foundation for learning various conditional controls. In this architecture, a trainable copy of the pre-trained model is linked to the original, fixed model through zero convolution layers, with initial weights set to zero. This design allows the trainable model to gradually incorporate new features during training while preventing the introduction of disruptive noise into the deep features of the large diffusion model. It is used in many multi-modal video generation strategies like  [36],  [148] [123],  [151]. In  [55], authors trained video generation module, Layout2Vid build of ModelScopeT2V  [125] with additional metadata about layouts, entities, and background. They merged bounding boxes, text, and image embedding with MLP. They enhanced ModelScopeT2V to consume additional conditional inputs from grounding tokens from layouts, textual entity description, and images to modulate the visual latent representation. LLM Grounded VDM  [52] added metadata with textual prompts termed dynamic scene layouts(DSL), but they took a training-free approach. FlowZero  [62] uses LLM to understand complex spatiotemporal dynamics from text. It generates a dynamic scene index containing scene descriptions, object layouts, and background motion patterns
Dynamic scene layout increases control and detail alignment between text and video but does not provide aesthetic infusion. Images can guide aesthetics if generated by a high-quality text-to-image diffusion model or provided as a reference. We will discuss a few papers incorporating images as a guidance mechanism.

### v2

@@ 3.2.1. Large Language Model As Director
The LLM-as-Director paradigm  [11] [97] [94] [98] revolutionizes video generation by employing a two-stage process: (1) an LLM Planner creates detailed narrative blueprints (keyframes, layouts, actions) and (2) a Video Generator Backbone produces intermediate frames (Fig. 9). This framework supports both zero-shot (training-free) and training-based approaches.
Free-Bloom  [21] demonstrates zero-shot capability through innovative techniques like joint noise sampling and DDIM-based dual-path interpolation  [99], though its LLM scripting potential is underutilized. VideoDirectorGPT  [11] enhances this with GPT-4’s comprehensive planning (layouts, bounding boxes) executed via Layout2Vid  [97]. The training-based LLM-grounded model  [100] further improves realism by learning spatiotemporal dynamics from the text.
Figure 8. The LLM as director approach utilizes LLM as the spatiotemporal director of the script, along with a separate video generation module that can understand the DSL (metadata) generated by LLM.  [67]
@@
Like VideoDirectorGPT, FlowZero  [22] adopted a zero-shot (training-free) approach; however, the LLM plays a more detailed role than Video Director GPT. It generates a detailed dynamic scene syntax (DSS), including scene descriptions, object arrangements, and background motion patterns. The DSS components direct an image diffusion model to generate videos with smooth object movements and consistent frame transitions. The theme of using dynamic scene layout is illustrated in Fig. 8. The LLM as director approach has limitations, as it is a two-stage architecture, and adding speech or integrating short clips will require modifications in the pipeline. We can extend the LLM-based divide-and-conquer approach by incorporating more specialized components, such as a model for generating reference images, a module for video creation, and a plug-and-play module for adding transitional clips and speech. That is achieved using the multi-agent framework, as described in the section. 3.2.2.
@@ 3.2.2. Multi stage/Agent Based Divide And Conquer Approach
Figure 10. Mora utilizes a multi-agent framework. The prompt selection agent enhances prompts with detailed instructions, the text-to-image agent generates images from input prompts, the image-to-image agent improves the quality of photos, the text-to-video agent generates video segments, and the video transition agent integrates these videos into a longer video  [93].
The LLM-as-Director paradigm  [11] [97] [94] [98] employs a two-stage process for video generation: (1) an LLM Planner creates narrative blueprints (keyframes, layouts, actions) and (2) a Video Generator Backbone produces intermediate frames (Fig. 9). This framework supports both zero-shot approaches like Free-Bloom  [21], which uses joint noise sampling and DDIM interpolation  [99], and training-based methods like the LLM-grounded model  [100]. VideoDirectorGPT  [11] enhances zero-shot generation through GPT-4’s detailed planning executed via Layout2Vid  [97], while addressing Free-Bloom’s limited script utilization.
@@ 3.2.3. Divide and Conquer Compositional/Transition Approach
Early long-video generation methods stitched short clips from standard models (diffusion/autoregressive) but struggled with transitions. SEINE  [25] innovated by framing transitions as masked diffusion, jointly denoising overlapping segments conditioned on boundary frames—though still requiring independent segment generation. Subsequent work improved continuity: MEVG  [101] anchored new clips to the final frames of predecessors, while MAVIN  [24] formalized transition learning as ”video infilling” by training on corrupted intermediates. Encoder-Empowered GAN  [102] enforced temporal coherence via recall mechanisms but sacrificed dynamic content flexibility.
These incremental advances culminated in VideoMerge [103], which reimagined the paradigm entirely. Instead of post hoc stitching, it preemptively ensures coherence through (1) adaptive noise blending to unify short and long temporal scales, (2) latent fusion for boundary-free transitions, and (3) prompt refinement for persistent identity. By addressing consistency at noise, latent, and semantic levels, VideoMerge achieves what prior segment-and-merge methods could not: holistic long-video synthesis without retraining, marking a shift from compositional fixes to native long-form generation.
The autoregressive approach ensures smooth transitions between frames by generating each frame conditioned on the previous ones, but its sequential nature makes it inherently slow for long video generation. The divide-and-conquer approach (see Section 3.2.1) can generate keyframes in parallel but faces challenges involving interpolation between frames with smooth transitions and achieving higher video quality while maintaining semantic consistency. Implicit generation 3.3 approaches combine the best of both worlds by generating complete videos directly from a model conditioned on user input without the need for interpolation (divide and conquer) or extrapolation (autoregressive) between frames. A summary of key papers exploring themes related to the Divide and Conquer approach is provided in Table  LABEL:table:divideandconquercatalog.
@@
SEINE [25] Long video with smooth transitions from short videos. Oct 2023 integration
@@
VideoMerge  [103] training free, merges short clips generated by pretrained text-to-video models March 2025 multi-stage
3.3. Implicit Video Generation using compressed Latent space
Figure 11. Using a Causal 3D VAE, Hunyuan Video compresses data into latent space. LLM-encoded text conditions Gaussian noise inputs, generating latents decoded into images/videos via the VAE decoder [105].
Implicit video generation synthesizes complete videos simultaneously through compact latent representations, employing spacetime compression, enhanced attention mechanisms, and hierarchical denoising  [20]. Unlike sequential approaches, models like Sora  [20] process entire videos via: (1) spacetime compression to latent patches (visualized in Fig 12), (2) ViT-based denoising, and (3) LLM-augmented CLIP-like conditioning. Open-source advances include FreeNoise  [106] for tuning-free semantic preservation, and GLOBER  [107] for efficient latent reconstruction. Hunyuan-Video  [105] (Fig. 11) advances the field through a diffusion-VAE hybrid architecture in compressed latent space, enabling both quality and long-form coherence—demonstrating implicit generation’s unique temporal synthesis capabilities from Sora to multimodal implementations.
Latent-space transformer-based video generation has progressed through key architectural innovations, beginning with Goku  [108]’s flow-based transformers for efficient joint image-video learning. Subsequent advances include REDUCIO!  [109]’s 3D VAE compression (64× more efficient than 2D VAEs) and Mochi 1  [110]’s 10B-parameter Asymmetric Diffusion Transformer for improved coherence. Meta’s Movie Gen![111] is a unified foundation model that generates high-quality images and videos from text prompts using efficient joint training in compressed latent space. The field’s current pinnacle is Cosmos  [84]’ World Foundation Model with 128× latent compression and two-phase training. However, persistent challenges in motion consistency and semantic alignment remain, evidenced by SORA’s  [1] one-minute generation limit and documented artifacts  [112]. Table 3 compares models using compressed latent spaces for video generation.
@@ 3.3. Implicit Video Generation using compressed Latent space
Mochi 1  [110] Asymmetric Diffusion Transformer (AsymmDiT) design October 2024
Hunyuan-Video  [105] Open source, MLLM Text Encoder and multiple video resolution . Dec 2024
Goku  [108] Flow based transformer and Vector-quantized variational autoencoder. Feb 2025
REDUCIO!  [109] Radical latent space compression, 3D VAE that compresses videos into ultra-compact motion representations. Nov 2024
Cosmos  [113] Physical AI based on world foundation models March 2025
Figure 13. An image is split into fixed-size patches, linearly embedded, augmented with position embeddings, and then passed through a Transformer encoder [57]
Beyond generation strategies, long video modeling requires effective tokenization - representing videos as compact units for efficient processing, as we discuss in the following section.
4. Long Video: Tokenization strategies
Long video tokenization strategies employ frame-level 4.1, temporal (3D Conv/VQ-VAE) 4.2, and hierarchical 4.3 approaches to efficiently encode spatiotemporal information while balancing computational demands and representation fidelity - critical for both generative and discriminative tasks, as detailed below.
4.1. Frame-Level Tokenization
Frame-level tokenization extends Vision Transformers [57] to video via fixed-size patches (e.g., 16×16 pixels) projected into latent tokens (Fig 13). Models like ViViT and MAGVIT [53] enhance this with spatiotemporal tokenization, hierarchical processing, and 3D embeddings. While effective, challenges remain in terms of attention complexity and motion-spatial trade-offs, driving the development of compressed token and hybrid architectures.
4.2. Temporal Tokenization (3D Conv/VQ-VAE)
Temporal tokenization compresses video into motion-aware latent representations, combining 3D convolutions with VQ-VAEs (Fig 14[46]. Pioneered by VideoGPT [47] and advanced in Hunyuan-Video [105] (Fig. 11), This approach tokenizes multiple spatiotemporal frames into a single compressed representation, such as MAGVIT [53] and cascaded approaches [114], address the fidelity-efficiency tradeoffs in long-video synthesis.
Figure 14. The proposed 3D-VQ architecture extends the 2D VQGAN [48] framework to model temporal dynamics. The system employs cascaded residual blocks in both encoder and decoder, combining 2D and 3D convolutions.  [115]
4.3. Hierarchical Tokenization
Hierarchical tokenization revolutionizes video compression by decomposing sequences into two optimized components: (1) high-fidelity keyframes that preserve critical visual details at sparse intervals and (2) parameterized motion representations that efficiently encode spatiotemporal dynamics between frames using compact neural descriptors. This asymmetric approach—inspired by human vision—achieves superior compression ratios while maintaining perceptual quality by decoupling spatial and temporal redundancies through learned priors, as proven in foundational work by HiTVideo [116]. Recent advancements integrate attention mechanisms for adaptive keyframe selection and diffusion-based motion estimation, further enhancing the framework’s efficiency and fidelity.
In addition to the tokenization strategy, another theme in the long videos is the use of input control mechanisms, such as text, bounding boxes, and images, for video guidance. We will discuss that in the next section.
@@ 5. Long Video: Input Control
Popular mechanisms for input conditioning of long videos include User Textual Prompt, User Textual Prompt with Scene Layout, and Image Input with Textual Prompt and Scene Layout, which we will discuss next.
@@ 5.1. User Textual Prompt
Text prompts serve as the primary conditioning method for video generation models, with implementations ranging from single-prompt to multi-prompt approaches. Early transformer-based models, such as DALL-E  [49] and CogVideo  [10], utilized autoregressive transformers on joint text-image token distributions, albeit with limitations to single prompts. Phenaki  [63] advanced this by incorporating T5X embeddings  [64] for sequential prompt conditioning, though facing coherence challenges in transitions. Contemporary solutions address these limitations through various approaches: Free-Bloom  [21] employs LLM-generated coherent prompts with spatial-temporal attention; LLM-Grounded Video Diffusion  [67] alternates between language guidance and denoising steps; VideoStudio  [119] modifies cross-attention mechanisms; and DirecT2V  [117] utilizes GPT-4 for step-by-step prompt generation. Fig. 16 illustrates the DirecT2V architecture [117], which modifies the attention block of U-Net and incorporates modulated self-attention. Text-only prompts can guide long video generation, but they lack the semantics necessary for fine-grained control over this process. In addition to frame descriptions, adding metadata, such as bounding boxes for entities like persons and cars, as well as background information, can help generate a more accurate depiction of videos and facilitate fine alignment between text and video.
@@ 5.2. User Textual Prompt with Scenes layout
Multi-modal input control mechanisms extend beyond text prompts to include images, edge maps, bounding boxes, and music, thereby enhancing generation with metadata such as entity trajectories and layouts (Fig. 8). Stable Diffusion  [45] pioneered this via cross-attention in UNet backbones (Fig. 15), while ControlNet  [120] introduced a trainable copy linked via zero convolutions to preserve pre-trained features. This architecture enables diverse applications  [121, 122, 123, 124]. Further advances include Layout2Vid in  [11] (extending ModelScopeT2V  [97] with layout/entity metadata), LLM-Grounded VDM’s training-free dynamic scene layouts  [67], and FlowZero’s spatiotemporal scene indices  [22]. Dynamic scene layout increases control and detail alignment between text and video but does not provide aesthetic infusion. Images can guide aesthetics if generated by a high-quality text-to-image diffusion model or provided as a reference. We will discuss a few papers incorporating images as a guidance mechanism.